### v1

@@ Abstract
Pathology image segmentation plays a pivotal role in computational pathology, which enables quantitative analysis of histological features for cancer diagnosis and prognosis. However, current segmentation methods encounter significant challenges in clinical applications, primarily due to the scarcity of high-quality, large-scale annotated pathology data and the constraints of fixed, narrowly defined object categories. To address these issues, this work aims to develop a segmentation foundation model capable of segmenting anything in pathology images using natural language. First, we establish PathSeg, the largest and most comprehensive dataset for pathology image semantic segmentation, derived from 17 publicly available datasets and comprising 275k image-mask-label triples. Our PathSeg dataset features a wide variety of 160 segmentation categories organized in a three-level hierarchy that covers 20 anatomical regions, 3 histological structures, and 61 object types. Next, we introduce PathSegmentor, the first text-prompted foundation model tailored for pathology image segmentation. With PathSegmentor, users can achieve semantic segmentation simply by providing a descriptive text prompt for the target category, thus eliminating the need to laboriously provide numerous spatial prompts like boxes or points for each instance. Extensive experiments on both internal and external datasets demonstrate the superior segmentation performance of PathSegmentor. It outperforms the group of specialized models, effectively handling a broader range of segmentation categories while maintaining a more compact model size. As a segmentation foundation model, PathSegmentor significantly surpasses other spatial-prompted and text-prompted models by 0.145 and 0.429 improvements in overall Dice scores, respectively, showcasing its remarkable robustness in segmenting complex objects and its effective generalization ability on external evaluations. Furthermore, we demonstrate that PathSegmentor’s versatile segmentation capabilities can effectively enhance the explainability of classification models for cancer diagnosis through feature importance estimation and imaging biomarker discovery. These interpretable outputs provide pathologists with evidence-based decision support, ultimately advancing precision oncology in clinical practice.
@@ 1 Introduction
The segmentation and identification of various structures in pathology images, including tissues, cells, and nuclei, represent a critical foundation for precision medicine [1, 2]. These analyses enable quantitative assessments, such as the calculation of cellular morphology including size, shape, texture, and other imaging features [3]. Recent breakthroughs in deep learning have revolutionized the field of pathology image segmentation, leading to the development of numerous specialized models [4, 5, 6, 7, 8, 9, 10] designed for particular tasks and datasets. Although these models often achieve high accuracy, their dependence on predefined segmentation categories limits their effectiveness in real-world clinical settings, which are characterized by diverse and dynamic conditions.
The emergence of segmentation foundation models presents a transformative solution to address these challenges. Segment Anything Model (SAM) [11] pioneered a novel prompt-driven paradigm that has since been expanded in subsequent segmentation foundation models [12, 13, 14]. These models enable flexible, interactive segmentation using various prompt types, including spatial prompts (e.g., points, boxes, scribbles), text prompts, reference prompts, or combinations thereof, significantly enhancing their adaptability across diverse applications. This versatility has generated substantial interest in medical imaging analysis. Several research efforts have explored the medical adaptation of segmentation foundation models [15, 16]. In multimodal image segmentation, models such as MedSAM [17], SAM-Med2D [18], and BiomedParse [19] have demonstrated effectiveness through distinct prompting strategies, including point/box prompts and text prompts. For specialized imaging modalities, researchers have developed dedicated solutions including $\mu$SAM [20] for microscopy analysis, SegVol [21] and SAT [22] for 3D radiology imaging, and SurgicalSAM [23] for endoscopic applications. Unlike traditional specialized segmentation models, foundation models not only handle a broader range of segmentation tasks, removing the need to train individual models for each application, but also exhibit enhanced generalization capabilities, enabling them to adapt effectively to diverse data distributions in real-world clinical settings.
@@
1) For dataset construction, we curate 275k image-mask-label triples from 17 publicly available pathology image segmentation datasets. Reflecting the inherent complexity of pathology, we reorganize semantic labels hierarchically, yielding 160 categories that encompass 20 anatomical regions, 3 histological structures, and 61 object types. To the best of our knowledge, this dataset, termed PathSeg, stands as the largest and most comprehensive benchmark for semantic segmentation of pathology images.
@@
Figure 1: Overview of PathSeg dataset. a, Treemap chart displaying the mask distribution across 17 pathology segmentation datasets in PathSeg. b, PathSeg dataset contains a total of 275k image-mask-label triples, significantly surpassing the amount of pathology data in other medical image segmentation benchmarks. c, Label distribution in PathSeg dataset illustrated by a Sankey diagram, highlighting the relationships among different anatomical regions, histological structures, and object types. The number in parentheses indicates the count of categories included in each grouping. d, Mask distribution in PathSeg dataset represented by pie charts, showing the number of masks across various anatomical regions and histological structures.
@@ 2.1 Overview of PathSeg Dataset
Extensive data scale. We construct the PathSeg dataset by collecting samples from 17 publicly available datasets (Fig. 1a) for pathology image semantic segmentation, all containing annotated pixel-level masks with categorical labels. The complete list of datasets including statistic information and download links is provided in Supplementary Table 1. The inclusion criteria comprises H&E-stained histopathology images from human tissue specimens and high-quality segmentation masks with precisely delineated contours. As a result, our PathSeg dataset comprises a total number of 275k image-mask-label triples (Fig. 1b), which significantly surpasses the pathology subsets of existing medical segmentation datasets such as MedSAM (9k) [17], SAM-Med2D-20M (1k) [18], and BiomedParseData (15k) [19]. The extensive data and semantic labels in PathSeg dataset provide strong support for the development of effective segmentation foundation models for pathology images.
@@
[…]
@@
In addition, Fig. 1c quantitatively illustrates the hierarchical label distributions, where the breast (47 labels), colon (17), lung (7), and prostate (7) contain the most diverse object types. Fig. 1d presents the mask distributions across anatomical regions and histological structures, respectively. Among anatomical regions, the prostate (24.7%), colon (13.6%), breast (7.1%), and lung (2.2%) exhibit the predominant proportions of masks. For histological structures, masks are distributed among tissue (55.9%), cell (28.7%), and nuclei (15.4%). The complete numbers of labels, images, and masks for each anatomical region, histological structure, and object type are provided in Supplementary Tables 2, 3, and 4, respectively.
@@ 2.2 Overview of PathSegmentor
Joint feature interaction module. As one of the key components in PathSegmentor, the joint feature interaction module integrates cross-attention layers, self-attention layers, and feed-forward networks (Fig. 2b and Methods). We create a set of learnable queries, along with image and text features as input to this module. Learnable queries first interact with image features through cross-attention, serving as adaptive filters to aggregate critical visual context. The vision-enhanced queries are then concatenated with text features and processed via self-attention, facilitating deep interactions where textual semantics guide the refinement of visual features. The final output from feed-forward networks is used to generate mask embeddings and class embeddings.
@@ 2.3 Segmentation Performance on PathSeg Dataset
Figure 3: Internal validation between PathSegmentor and specialized models across 16 datasets from PathSeg dataset. a, Bar charts of Dice scores on each dataset, calculated using bootstrapping with 95% confidence interval. Statistical significance is indicated as **$p\leq 0.01$, *$p\leq 0.05$. P-values are computed using one-sided, paired Student’s $t$-tests. b, Scatter plots of overall performance for each model with the number of models, number of categories, and total model size.
@@ 2.3.1 Comparison with Specialized Segmentation Models
First, we compare PathSegmentor against three state-of-the-art specialized segmentation models (nnU-Net, DeepLabV3+, SAM-Path) on internal datasets. Each specialized model is trained using dataset-optimized configurations through their official implementations. The quantitative results on 16 datasets are shown in Fig. 3a and Supplementary Table 6. Among the three specialized models, nnU-Net achieves superior Dice scores on 11 out of 16 datasets, while SAM-Path and DeepLabV3+ excel on 3 and 2 datasets, respectively. Notably, as a segmentation foundation model, PathSegmentor further outperforms all three on 14 out of 16 datasets with statistical significance ($p\leq 0.05$, one-sided, paired Student’s $t$-tests). Specifically, on SegPath, a challenging multi-category dataset containing eight cell types, PathSegmentor scores 0.442, achieving a 0.28 performance improvement over DeepLabV3+’s score of 0.162. The superiority of PathSegmentor can be attributed to large-scale training data and hierarchical semantic modeling. Our constructed PathSeg dataset facilitates robust feature learning across diverse pathology patterns, whereas specialized models are constrained by the sizes of individual datasets. Additionally, hierarchical textual inputs explicitly encode structured pathology knowledge to enhance morphological discrimination. These results illustrate how foundation models trained on comprehensive data with semantic priors can overcome the limitations of traditional paradigms that rely on one specialized model per dataset.
We also compare the overall performance of four models by averaging the Dice scores of all samples on 16 internal datasets. As illustrated in Fig. 3b, PathSegmentor achieves a superior overall Dice score of 0.671, compared to 0.502 for nnU-Net, 0.462 for DeepLabV3+, and 0.472 for SAM-Path. PathSegmentor exhibits three key advantages. 1) Unified architecture: a single PathSegmentor model outperforms a group of 16 specialized nnU-Net models with a large margin of 0.169 Dice score. 2) Category scalability: PathSegmentor can support all 160 semantic classes in the PathSeg dataset, which is 40$\times$ more than the average capacity of specialized models, each handling approximately 4 classes based on the total of 16 models for 61 object types. 3) Model efficiency: PathSegmentor has a significantly smaller model size of 0.45B parameters and achieves a 75% size reduction compared to the group of 16 SAM-Path models, which have 1.86B parameters. This makes PathSegmentor more suitable for clinical deployment. In addition, we present more detailed anatomical region level results in Supplementary Table 12, histological structure level results in Supplementary Table 13, and object type level results in Supplementary Table LABEL:tab:performance_across_targets.
@@ 2.3.2 Comparison with Spatial-prompted Segmentation Foundation Models
Figure 4: Internal validation between PathSegmentor and spatial-prompted segmentation foundation models across 16 datasets from PathSeg dataset. Bar charts of Dice scores for overall performance and each dataset, calculated using bootstrapping with 95% confidence interval. Statistical significance is indicated as **$p\leq 0.01$. P-values are computed using one-sided, paired Student’s $t$-tests.
@@
1) Irregular object shape. Pathological objects often exhibit irregular morphologies due to lesions, inflammation, or tumor infiltration. To quantify irregularity, we employ the metric derived from circularity [33] (Methods). Fig. 5a and Supplementary Table 8 demonstrate that while both MedSAM and PathSegmentor attain Dice scores exceeding 0.9 for regular-shaped objects, their performance diverges significantly with increasing shape irregularity. MedSAM exhibits a parabolic decline in Dice score that drops to 0.282, whereas PathSegmentor sustains stable performance at 0.790 with only marginal degradation. This performance gap under maximal irregularity highlights PathSegmentor’s superior shape robustness. Visual confirmation is provided in Fig. 5b, where PathSegmentor precisely delineates the complex boundaries of irregular tissues across different anatomical regions that challenge the spatial-prompted approaches.
@@
3) High instance density. Pathology images frequently exhibit exceptionally high instance densities, posing significant challenges for segmentation tasks. We use the number of distinct instances within each object mask to indicate instance density. In Fig. 5e, statistic analysis of our PathSeg test set reveals an instance count distribution: 90% of masks contain 0$\sim$72 instances, while the upper 10% contain cases extending beyond this range, reaching a maximum count surpassing 800 instances within a mask. This distribution underscores the inherent high density in pathology images. Fig. 5f and Supplementary Table 10 illustrate the variation in segmentation results of PathSegmentor and MedSAM as a function of instance density. When instance density rises, MedSAM suffers a more than 0.3 Dice decline, which is attributed to the union box prompts lacking sufficient spatial information to resolve individual instances in crowded regions. In contrast, PathSegmentor maintains robust performance with a Dice score around 0.7, demonstrating superior capability in segmenting objects with dense instances through text prompts.
4) Prompt efficiency analysis. Furthermore, we analyze prompt efficiency by evaluating the trade-off between performance and the number of prompts across different prompting strategies. This is clinically important as the required number of prompts directly correlates with pathologist annotation time. Fig. 5g and Supplementary Table 11 compare four segmentation approaches in terms of their performance and the number of prompts required per mask, including PathSegmentor, MedSAM (Union Box), MedSAM (Instance Box), and Instance Box which directly uses the instance boxes as prediction. PathSegmentor demonstrates superior single-prompt efficiency, achieving a 0.145 higher Dice score than MedSAM (Union Box) when limited to one prompt per image. The enhanced performance of MedSAM (Instance Box) and Instance Box primarily results from their extensive reliance on precise spatial localization cues. This requires approximately 15$\times$ more prompts per mask overall, specifically 12 for tissues, 16 for nuclei, and 28 for cells. However, this increased demand creates significant challenges for clinical workflows that limit practical adoption. Fig. 5h visually demonstrates these challenges in extreme cases, where single patches already contains 288 nuclei and 396 cells. When extended to whole-slide images that typically comprise thousands of such patches, the annotation complexity becomes prohibitively expensive, potentially requiring annotation of millions of individual objects. These results demonstrate PathSegmentor’s advantage in balancing accuracy with practical usability through its efficient text-prompted paradigm for real-world pathology practice.
@@ 2.3.3 Comparison with Text-prompted Segmentation Foundation Models
Fig. 6b presents the comparison results between PathSegmentor and BiomedParse across three specific levels, including anatomical region, histological structure, and object type. For anatomical regions, we focus on five primary areas with over 1,000 masks, including prostate, colon, breast, lung, and unspecified, whereas the results for all regions can be found in Supplementary Table 12. The results indicate that PathSegmentor demonstrates a significant advantage in segmentation accuracy across these key anatomical regions. In terms of histological structure, PathSegmentor significantly outperforms BiomedParse across tissue, cell, and nuclei, ranging from large and irregular features to small and dispersed ones (Supplementary Table 13). Regarding object type, we showcase the top eight objects with the highest number of masks, including epithelial, smooth muscle cells, prostate healthy glands, prostate cancer, leukocytes, red blood cells, prostate tumor, and lymphocyte, which have mask counts ranging from 15k to 30k. The results for all object types are provided in Supplementary Table LABEL:tab:performance_across_targets. This highlights PathSegmentor’s ability to segment a diverse array of biological entities effectively. The segmentation results of all 160 categories in PathSeg dataset are presented in Supplementary Table LABEL:tab:performance_across_categories.
@@ 2.4 External Evaluation
We assess the generalization performance of PathSegmentor and other segmentation models across pathology images acquired from diverse clinical environments. The external validation employs datasets that are completely independent from the internal training datasets, while ensuring that the evaluated categories were previously learned during model training. In the PathSeg dataset, categories are organized into three hierarchical levels, which presents a practical challenge: few external datasets contain categories that perfectly matched all three levels of our internal categories. To address this issue while maintaining scientific rigor, we define a category as valid for external evaluation if it shares the same [object type] with any category present in internal datasets. The CoNSeP dataset [8] meets this criterion, consisting of 121 samples across 6 cell types from lung regions.
@@
As segmentation foundation models, PathSegmentor, MedSAM, SAM-Med2D, and BiomedParse are directly tested on the external dataset. For specialized models, given their tailored configurations for each dataset, we systematically evaluate the specialized models that can handle the corresponding internal category for each external category (Supplementary Table 16). For example, to evaluate the performance on colon-cell-malignant epithelial in CoNSeP, we adopt the specialized models trained on WSSS4LUAD, which involves the category of tumor epithelial tissue. For the categories appearing in multiple datasets, including colon-cell-healthy epithelial, colon-cell-fibroblast, and colon-cell-endothelial, we adopt the results of the best-performing specialized models. The external validation results are shown in Fig. 7 and Supplementary Table 17.
@@
Figure 7: External evaluation of PathSegmentor compared to foundation and specialized models on the CoNSeP dataset. a, Overall Dice scores with bootstrapped 95% confidence interval. Statistical significance is indicated as **$p\leq 0.01$. b, Per-category Dice scores for all object types. PathSegmentor results are visually emphasized through enlarged markers and annotated with rank positions. c, A scatter plot illustrating the number of instances and the maximum inter-instance distance per image, with horizontal and vertical error bars.
@@
The comparison results of overall performance in Fig. 7a reveal three key findings. First, foundation models (PathSegmentor, MedSAM, SAM-Med2D, and BiomedParse) collectively demonstrate significantly stronger performance than specialized models (nnU-Net, DeepLabV3+, and SAM-Path), highlighting their superior generalization capability in real-world clinical settings. Second, within the group of foundation models, the text-prompted model (BiomedParse) outperforms both spatial-prompted models (MedSAM and SAM-Med2D) by substantial margins of 0.047 and 0.069 in Dice scores, respectively. Finally, PathSegmentor establishes state-of-the-art performance, further surpassing BiomedParse by an additional 0.074 Dice score, which clearly shows its enhanced effectiveness.
Specifically, we focus on the performance of each category in the external dataset, all of which are colon cells. Fig. 7b illustrates that PathSegmentor with text prompts consistently achieves superior performance for the five cells (malignant epithelial, healthy epithelial, muscle, inflammatory, and fibroblast cells), outperforming all competing models. Particularly, it reaches Dice scores of 0.624 for malignant epithelial and 0.596 for healthy epithelial cells, representing 0.162 and 0.331 improvements over MedSAM (Union Box), the best spatial-prompted model. However, we notice an important exception. Although text-prompted models excel for most cell types, both PathSegmentor and BiomedParse underperform spatial-prompted models (MedSAM and SAM-Med2D) on endothelial cells. This divergence motivates our investigation of cellular distribution patterns.
@@
To quantify these differences, we visualize each sample’s amount-dispersion distribution in Fig. 7c, where amount represents the instance count per category in each image, and dispersion measures the maximum inter-instance distance within each category mask. Distribution statistics of mean and standard deviation for each category are detailed in Supplementary Table 18. The analysis reveals that the five better-performance cell types exhibit both abundant instances and wide spatial distributions, whereas endothelial cells appear sparsely with clustered arrangements. This explains PathSegmentor’s strength in segmenting high-density, high-dispersion cellular regions, while spatial prompts provide more precise localization cues for few and concentrated cells.

### v2

@@ Abstract
Pathology image segmentation plays a pivotal role in computational pathology, which enables quantitative analysis of histological features for cancer diagnosis and prognosis. However, current segmentation methods encounter significant challenges in clinical applications, primarily due to the scarcity of high-quality, large-scale annotated pathology data and the constraints of fixed, narrowly defined object categories. To address these issues, this work aims to develop a segmentation foundation model capable of segmenting anything in pathology images using natural language. First, we establish PathSeg, the largest and most comprehensive dataset for pathology image semantic segmentation, derived from 21 publicly available datasets and comprising 275k image-mask-label triples. Our PathSeg dataset features a wide variety of 160 segmentation categories organized in a three-level hierarchy that covers 20 anatomical regions, 3 histological structures, and 61 object types. Next, we introduce PathSegmentor, a text-prompted foundation model tailored for pathology image segmentation. With PathSegmentor, users can achieve semantic segmentation simply by providing a descriptive text prompt for the target category, thus eliminating the need to laboriously provide numerous spatial prompts like boxes or points for each instance. Extensive experiments on both internal and external datasets demonstrate the superior segmentation performance of PathSegmentor. It outperforms the group of specialized models, effectively handling a broader range of segmentation categories while maintaining a more compact model size. As a segmentation foundation model, PathSegmentor significantly surpasses other spatial-prompted and text-prompted models by 0.145 and 0.429 improvements in overall Dice scores, respectively, showcasing its remarkable robustness in segmenting complex objects and its effective generalization ability on external evaluations. Furthermore, we demonstrate that PathSegmentor’s versatile segmentation capabilities can effectively enhance the explainability of classification models for cancer diagnosis through feature importance estimation and imaging biomarker discovery. These interpretable outputs provide pathologists with evidence-based decision support, ultimately advancing precision oncology in clinical practice.
@@ 1 Introduction
The segmentation and identification of various structures in pathology images, including tissues, cells, and nuclei, represent a critical foundation for precision medicine [1, 2]. These analyses enable quantitative assessments, such as the calculation of cellular morphology including size, shape, texture, and other imaging features [3]. Recent breakthroughs in deep learning have revolutionized the field of pathology image segmentation, leading to the development of numerous specialized models [4, 5, 6, 7, 8, 9, 10, 11, 12] designed for particular tasks and datasets. Although these models often achieve high accuracy, their dependence on predefined segmentation categories limits their effectiveness in real-world clinical settings, which are characterized by diverse and dynamic conditions.
[…]
@@
1) For dataset construction, we curate 275k image-mask-label triples from 21 publicly available pathology image segmentation datasets. Reflecting the inherent complexity of pathology, we reorganize semantic labels hierarchically, yielding 160 categories that encompass 20 anatomical regions, 3 histological structures, and 61 object types. To the best of our knowledge, this dataset, termed PathSeg, stands as the largest and most comprehensive benchmark for semantic segmentation of pathology images.
@@
Figure 1: Overview of PathSeg dataset. a, Treemap chart of 21 pathology segmentation datasets in PathSeg (internal, blue; external, red), where area size and color intensity indicate the number of masks and labels, respectively. b, PathSeg contains a total of 275k image-mask-label triples, significantly surpassing the amount of pathology data in other medical image segmentation benchmarks. c, Label distribution in PathSeg illustrated by a Sankey diagram, highlighting the relationships among different anatomical regions (AR), histological structures (HS), and object types (OT). The number in parentheses indicates the count of categories included in each grouping. d, Donut charts visualizing mask distribution across three hierarchical levels in PathSeg.
@@ 2.1 Overview of PathSeg Dataset
Extensive data scale. We construct the PathSeg dataset by collecting samples from 21 publicly available datasets (Fig. 1a) for pathology image semantic segmentation, all containing annotated pixel-level masks with categorical labels. The complete list of datasets including statistic information and download links is provided in Supplementary Table 1. The inclusion criteria comprises H&E-stained histopathology images from human tissue specimens and high-quality segmentation masks with precisely delineated contours. As a result, our PathSeg dataset comprises a total number of 275k image-mask-label triples (Fig. 1b), which significantly surpasses the pathology subsets of existing medical segmentation datasets such as BiomedParseData (15k) [21], MedSAM (9k) [19], and SAM-Med2D-20M (1k) [20]. The extensive data and semantic labels in PathSeg dataset provide strong support for the development of effective segmentation foundation models for pathology images.
@@
$\text{[anatomical region] - [histological structure] - [object type]},$
@@
In addition, Fig. 1c quantitatively illustrates the hierarchical label distributions, where the breast (47 labels), colon (17), lung (7), and prostate (7) contain the most diverse object types. Fig. 1d presents the mask distributions across three levels. Among anatomical regions, the prostate (24.7%), colon (13.6%), breast (7.1%), and lung (2.3%) exhibit the predominant proportions of masks. For histological structures, masks are distributed among tissue (55.8%), cell (28.7%), and nuclei (15.5%). We show the top 15 object types. The complete numbers of labels, images, and masks for each anatomical region, histological structure, and object type are provided in Supplementary Tables 2, 3, and 4, respectively.
@@ 2.2 Overview of PathSegmentor
Joint feature interaction module. As one of the key components in PathSegmentor, the joint feature interaction module integrates cross-attention layers, self-attention layers, and feed-forward networks (Fig. 2b and Section 4.2). We create a set of learnable queries, along with image and text features as input to this module. Learnable queries first interact with image features through cross-attention, serving as adaptive filters to aggregate critical visual context. The vision-enhanced queries are then concatenated with text features and processed via self-attention, facilitating deep interactions where textual semantics guide the refinement of visual features. The final output from feed-forward networks is used to generate mask embeddings and class embeddings.
@@ 2.3 Segmentation Performance on PathSeg Dataset
Figure 3: Internal validation between PathSegmentor and specialized models across 16 datasets from PathSeg dataset. a, Bar charts of Dice scores on each dataset, calculated using bootstrapping with 95% confidence interval. Statistical significance is indicated as $p\leq 0.01$, $p\leq 0.05$. P-values are computed using one-sided, paired Student’s $t$-tests. b, Scatter plots of overall performance for each model with the number of models, number of categories, and total model size.
@@ 2.3.1 Comparison with Specialized Segmentation Models
First, we compare PathSegmentor against three state-of-the-art specialized segmentation models (nnU-Net, DeepLabV3+, SAM-Path) on internal datasets. Each specialized model is trained using dataset-optimized configurations through their official implementations. The quantitative results on 16 datasets are shown in Fig. 3a and Supplementary Table 6. Among the three specialized models, nnU-Net achieves superior Dice scores on 11 out of 16 datasets, while SAM-Path and DeepLabV3+ excel on 3 and 2 datasets, respectively. Notably, as a segmentation foundation model, PathSegmentor further outperforms all three on 14 out of 16 datasets with statistical significance ($p\leq 0.05$, one-sided, paired Student’s $t$-tests). Specifically, on SegPath, a challenging multi-category dataset containing eight cell types, PathSegmentor scores 0.442, achieving a 0.280 performance improvement over DeepLabV3+’s score of 0.162. The superiority of PathSegmentor can be attributed to large-scale training data and hierarchical semantic modeling. Our constructed PathSeg dataset facilitates robust feature learning across diverse pathology patterns, whereas specialized models are constrained by the sizes of individual datasets. Additionally, hierarchical textual inputs explicitly encode structured pathology knowledge to enhance morphological discrimination. These results illustrate how foundation models trained on comprehensive data with semantic priors can overcome the limitations of traditional paradigms that rely on one specialized model per dataset.
[…]
@@ 2.3.2 Comparison with Spatial-prompted Segmentation Foundation Models
Figure 4: Internal validation between PathSegmentor and spatial-prompted segmentation foundation models across 16 datasets from PathSeg dataset. Bar charts of Dice scores for overall performance and each dataset, calculated using bootstrapping with 95% confidence interval. Statistical significance is indicated as $p\leq 0.01$. P-values are computed using one-sided, paired Student’s $t$-tests.
@@
1) Irregular object shape. Pathological objects often exhibit irregular morphologies due to lesions, inflammation, or tumor infiltration. To quantify irregularity, we employ the metric derived from circularity [35] (Section 4.5). Fig. 5a and Supplementary Table 8 demonstrate that while both MedSAM and PathSegmentor attain Dice scores exceeding 0.9 for regular-shaped objects, their performance diverges significantly with increasing shape irregularity. MedSAM exhibits a parabolic decline in Dice score that drops to 0.282, whereas PathSegmentor sustains stable performance at 0.790 with only marginal degradation. This performance gap under maximal irregularity highlights PathSegmentor’s superior shape robustness. Visual confirmation is provided in Fig. 5b, where PathSegmentor precisely delineates the complex boundaries of irregular tissues across different anatomical regions that challenge the spatial-prompted approaches.
@@
[…]
@@ 2.3.3 Comparison with Text-prompted Segmentation Foundation Models
Fig. 6b presents the comparison results between PathSegmentor and BiomedParse across three specific levels, including anatomical region, histological structure, and object type. For anatomical regions, we focus on five primary areas with over 1,000 masks, including prostate, colon, breast, lung, and unspecified, whereas the results for all regions can be found in Supplementary Table 12. The results indicate that PathSegmentor demonstrates a significant advantage in segmentation accuracy across these key anatomical regions. In terms of histological structure, PathSegmentor significantly outperforms BiomedParse across tissue, cell, and nuclei, ranging from large and irregular features to small and dispersed ones (Supplementary Table 13). Regarding object type, we showcase the top eight objects with the highest number of masks, including epithelial, smooth muscle cells, prostate healthy glands, prostate cancer, leukocytes, red blood cells, prostate tumor, and lymphocyte, which have mask counts ranging from 15k to 30k. The results for all object types are provided in Supplementary Table LABEL:tab_supp:performance_across_targets. This highlights PathSegmentor’s ability to segment a diverse array of biological entities effectively. The segmentation results of all 160 categories in PathSeg dataset are presented in Supplementary Table LABEL:tab_supp:performance_across_categories.
@@ 2.4 External Evaluation
We assess the generalization performance of PathSegmentor and other segmentation models across pathology images acquired from diverse clinical environments. The external validation employs datasets that are completely independent from the internal training datasets, while ensuring that the evaluated categories were previously learned during model training.
In our experiments, we conduct external evaluation using 5 datasets where the [object type] in each external category align with categories present in our internal datasets. Specifically, CPM17, CPM15, Kumar, and Lizard datasets contain the nuclei category, while CoNSeP dataset includes 6 types of colon cells.
@@
As segmentation foundation models, PathSegmentor, MedSAM, SAM-Med2D, and BiomedParse are directly tested on the external datasets. For specialized models, given their tailored configurations for each dataset, we systematically evaluate the specialized models that can handle the corresponding internal category for each external category. As illustrated in Supplementary Table 16, for example, to evaluate the performance on colon-cell-malignant epithelial in CoNSeP, we adopt the specialized models trained on WSSS4LUAD, which involves the category of tumor epithelial tissue. For the categories appearing in multiple training datasets, including colon-cell-healthy epithelial, colon-cell-fibroblast, and colon-cell-endothelial, we adopt the results of the best-performing specialized models. The external validation results are shown in Fig. 7 and Supplementary Table 17.
@@
Figure 7: External evaluation of PathSegmentor compared to foundation and specialized models. a-e, Overall Dice scores with bootstrapped 95% confidence interval on 5 external datasets. Statistical significance is indicated as $p\leq 0.01$. f, Per-category Dice scores for all object types on CoNSeP dataset. PathSegmentor results are visually emphasized through enlarged markers and annotated with rank positions. g, A scatter plot illustrating the number of instances and the maximum inter-instance distance per image on CoNSeP dataset, with horizontal and vertical error bars.
@@
The comparison results of overall performance in Fig. 7a-e reveal that foundation models (PathSegmentor, MedSAM, SAM-Med2D, and BiomedParse) collectively demonstrate significantly stronger performance than specialized models (nnU-Net, DeepLabV3+, and SAM-Path), highlighting their remarkable generalization capability in real-world clinical settings. Among these, PathSegmentor establishes consistent state-of-the-art performance across all external datasets. Specifically for nuclei segmentation, PathSegmentor achieves 0.705, 0.485, 0.427, and 0.319 Dice scores on CPM17, CPM15, Kumar, and Lizard, respectively, demonstrating superior performance in dense object segmentation tasks. On CoNSeP dataset, PathSegmentor surpasses the second best-performing model BiomedParse by 0.074 Dice score, which clearly shows its enhanced effectiveness.
@@
Furthermore, we focus on the performance of each category in CoNSeP dataset, all of which are colon cells. Fig. 7b illustrates that PathSegmentor with text prompts consistently achieves superior performance for the five cells (malignant epithelial, healthy epithelial, muscle, inflammatory, and fibroblast cells), outperforming all competing models. Particularly, it reaches Dice scores of 0.624 for malignant epithelial and 0.596 for healthy epithelial cells, representing 0.162 and 0.331 improvements over MedSAM (Union Box), the best spatial-prompted model. However, we notice an important exception. Although text-prompted models excel for most cell types, both PathSegmentor and BiomedParse underperform spatial-prompted models (MedSAM and SAM-Med2D) on endothelial cells. This divergence motivates our investigation of cellular distribution patterns. To quantify these differences, we visualize each sample’s amount-dispersion distribution in Fig. 7c, where amount represents the instance count per category in each image, and dispersion measures the maximum inter-instance distance within each category mask. Distribution statistics of mean and standard deviation for each category are detailed in Supplementary Table 18. The analysis reveals that the five better-performance cell types exhibit both abundant instances and wide spatial distributions, whereas endothelial cells appear sparsely with clustered arrangements. This explains PathSegmentor’s strength in segmenting high-density, high-dispersion cellular regions, while spatial prompts provide more precise localization cues for few and concentrated cells.